3. A spring loaded gun fires a bullet into a block of wood hanging from a frictionless pendulum. a. How fast was the bullet going when it left the gun?
Spring: k = 5,000 N/m
Bullet: m=20.0g
Block of wood: m= 1.00kg

b. What was the velocity of the block the moment after the bullet hit it?

c. How high would the wooden block swing if the bullet becomes imbedded in it?
1
Expert's answer
2020-06-15T10:34:20-0400

a)


"v=v_i"

b)


"(m+M)V=mv_i\\(0.02+1)V=0.02v_i\\V=0.0196v_i"

c)


"(m+M)gh=0.5mv_i^2\\(0.02+1)(9.8)h=0.02v_i^2\\h=0.002v_i^2"